d²V/dt² -4V = 8e^(2t)

When t=0 and V=4 and dV/dt = 6


1
Expert's answer
2020-07-16T18:20:00-0400

This is a second order differential equation with constant coefficients. The characteristic polynomial of the homogeneous equation is "r^2-4=0". The roots of this polynomials are "r_1=2 \text{ and } r_2=-2". The general solution of the homogeneous equation is then "V = c_1e^{2t}+c_2e^{-2t}".

A particular solution of the nonhomogeneous equation is "cte^{2t}", where "c" is a constant. To determine it, we replace "V=cte^{2t} \text{ and }\frac{d^2V}{dt^2}=4c(1+t)e^{2t}" in the equation, we obtain

"4c(1+t)e^{2t}-4cte^{2t}=8t\implies 4ce^{2t}=8e^{2t}\implies c=2."

The solution is then

"V(t) = c_1e^{2t}+c_2e^{-2t}+2te^{2t}."

To determine "c_1\text{ and }c_2" we use the initial conditions:

"V(0) = c_1+c_2=4" and

"\frac{dV}{dt}=(2c_1+4t+2)e^{2t}-2c_2e^{-2t}\Rightarrow\frac{dV}{dt}(0)=2c_1-2c_2+2=6\Rightarrow c_1-c_2=2".

Then "c_1=3\text{ and }c_2=1". We obtain finally

"V(t)=3e^{2t}+e^{-2t}+2te^{2t}."